# A portable LED-based diamond magnetometer for outreach and teaching labs

### v1

@@ Abstract
Hollis Williams1,2, Alex Newman1, Stuart Graham1, Colin Stephen1, Gavin Morley1
1Department of Physics, University of Warwick, Coventry CV4 7AL, United Kingdom
2Current address: Theoretical Sciences Visiting Program, Okinawa Institute of Science and Technology Graduate University, Onna 904-0495, Japan
Contact: gavin.morley@warwick.ac.uk, holliswilliams@hotmail.co.uk
1 Introduction
@@
In this Note, we present a portable NV magnetometer which uses a high-power green LED instead of a laser. This magnetometer prioritizes safety, portability, reliable performance, and direct visualization for outreach and undergraduate settings. Although it does not reach the affordability of the very low cost modular options which are available in the literature, the magnetometer which we present here is tailored for teaching and outreach environments where clarity of the fluorescence signal and ease and reliability of the demonstration are essential. The use of an LED reduces enables both the green excitation and red photoluminescence to be seen with the naked eye, which is especially effective in public demonstrations. Low power (few-milliwatt) 532 nm laser modules can also be used in teaching-lab implementations of ODMR experiments, LED illumination requires no laser safety procedures or alignment, making it particularly suitable for outreach demonstrations and environments with minimal supervision.
@@
Figure 1: (Left) Schematic of the experiment. The gray squares denote the two block magnets attached to the circuit board and the pink square denotes the NV diamond. (Right) Schematic of the prism and diamond placement (not to scale). The bottom rectangle denotes the circuit board and the rectangle between the prism and the circuit board is the NV diamond placed on top of a thin metal pad. The gap between the prism and the light mixing rod is exaggerated to indicate that a filter is placed in between.
Figure 2: Excerpt of the top copper Gerber layer of the printed circuit board showing the integrated microwave loop used for ODMR. The circular pad on the right connects to the SMA input. A straight copper trace routes the microwave signal to the printed single-turn loop in the middle of the board, which generates the oscillating magnetic field to drive the NV spin transitions. The diamond sits directly above this loop during operation. Full PCB design files are provided in the Supplementary Material.
Figure 3: Measured ODMR spectrum from the LED-excited diamond magnetometer, acquired using the same operating settings as the demonstration in Fig. 5. The dashed line marks the working point, defined as the frequency at which the ODMR amplitude exhibits its maximum slope $dV/df$.
Figure 4: Photo of the demonstration setup.
2 Concept and Demonstration
@@
The central idea of the device is to use a high-power green LED to excite a millimetre-scale diamond containing a high concentration of NV- centers, and to detect the resulting red photoluminescence (PL) while applying microwave radiation near the NV spin resonances. A simplified schematic of the optical and electronic layout is shown in Fig. 1(a). The LED output is fed into a short hexagonal light-mixing rod, which homogenizes the illumination and delivers it to a right-angle prism. The prism directs the green light onto the diamond mounted on the printed circuit board, whilst simultaneously collecting the red photoluminescence and returning it through the rod to a long-pass filter and amplified photodiode. Since the green excitation and red fluorescence are both easily visible to the naked eye, students can directly observe the physical processes responsible for the ODMR signal.
@@
To clarify the optical and mechanical interface on the circuit board, Fig. 1(b) provides a detailed view of the prism-rod-diamond assembly. The diamond is bonded to a reflective metal pad on the board, the prism is placed on top of the diamond, and the face of the mixing rod is aligned so that it presses against the face of the prism. This compact geometry provides efficient PL collection without the need for lenses and keeps alignment requirements minimal. The board incorporates a printed single-turn loop antenna which provides the oscillating magnetic field required for ODMR. The loop is etched into the top copper layer and is driven from the SMA connector through a short copper transmission line, as shown in Fig. 2. The diamond is positioned directly above the loop to ensure strong near-field coupling. An excerpt of the top copper layer is included in the figure for clarity, and the full Gerber files are available in the Supplementary Material.
Microwave excitation for ODMR is supplied by a compact module. In continuous-wave operation, the LED is held at constant intensity whilst the microwave frequency is swept. When the microwave field is resonant with the $\lvert m_{s}=0\rangle\!\to\!\lvert m_{s}=\pm 1\rangle$ transitions, the PL decreases, producing an ODMR dip in the fluorescence. With a bias magnetic field supplied by two small permanent magnets, the degeneracy of the $\lvert m_{s}=\pm 1\rangle$ states is lifted and a pair of resonances is observed. A representative ODMR spectrum obtained with the LED magnetometer is shown in Fig. 3. The spectrum was recorded using the same operating configuration as the magnetometer demonstration (identical LED driver setting, microwave source power, and Red Pitaya lock-in configuration), ensuring that it reflects the actual working conditions of the device. The demodulated X-channel exhibits the characteristic dispersive line shape with two resonances corresponding to different NV orientations under the applied bias field, whereas the Y-channel remains close to zero. We indicate the magnetometer working point by marking the frequency at which the ODMR amplitude has the largest slope $dV/df$, which maximizes magnetic field responsivity.
@@
The magnetometer has been used successfully in outreach activities. The entire setup packs into two small plastic boxes and can be reassembled in a few minutes (shown in Fig. 4). At a recent event in our Physics Department, students viewed the PL signal on a laptop whilst magnetic objects were brought near the sensor. For simple demonstrations, it is sufficient to observe the raw PL signal, which changes visibly in response to nearby magnetic objects. Fig. 5 shows the demodulated PL signal whilst a steel Allen key was moved close to the magnetometer. It can be seen that a spike forms as the object is brought near, giving a compelling demonstration that the device functions as a magnetometer.
@@
3 Conclusion
We have presented a portable NV center magnetometer which uses LED illumination instead of a laser, making it especially suitable for outreach and undergraduate laboratory settings. The device produces clear ODMR spectra and responds visibly to nearby magnetic objects, enabling both qualitative demonstrations and quantitative measurements. During outreach events the magnetometer proved easy to assemble and operate, with students able to observe its response in real time. In an instructional laboratory, the system can support introductory activities, as well as more advanced exercises involving magnet alignment, bias-field estimation, or parameter optimisation.
Author Declarations
The authors have no conflicts to disclose.
Acknowledgments
Alex Newman’s PhD studentship is funded by an EPSRC iCASE award to UKNNL (United Kingdom National Nuclear Laboratory). This work received funding from the National Nuclear Laboratory’s Science and Technology programme (Decontamination and Decommissioning Core Science). This work is also supported by Innovate UK grant 10003146, EPSRC grant EP/V056778/1 (Prosperity Partnership with Element Six), an EPSRC Impact Acceleration Account (IAA) award, the EPSRC Q-BIOMED Hub EP/Z533191/1, the EPSRC Quantum Computation and Simulation Hub EP/T001062/1, and STFC grant number ST/W006561/1. This research is funded in part by the Gordon and Betty Moore Foundation through Grant GBMF12328, DOI 10.37807/GBMF12328. This material is based upon work supported by the Alfred P. Sloan Foundation under Grant No. G-2023-21130.
@@ Acknowledgments
LED Magnetometer
Design
In this section, we discuss the components used to build the magnetometer (these are listed in Table I along with the model, estimated cost, and suggested supplier). The prices were checked on 22nd November 2025 and there may be some variation among different vendors. It is also possible to lower the cost by looking for used parts. The overall estimated cost of the device is around \$4500, similar to the setup described in [18]. In some cases, cheaper alternatives can be used to decrease the cost (for example, filtered glass could be used instead of a dichroic shortpass filter). We also expect that some re-design of the LED part could result in a substantial price reduction.
@@
Red Pitaya STEMLab 125-14 \$400 Red Pitaya
Microwave source TPI-1005-A \$300 RF Consultant
Dichroic 600 nm shortpass filter 69-180 \$100 Edmund Optics
Amplified photodetector PDA100A2 \$400 Thorlabs
T-Cube LED driver LEDD1B \$300 Thorlabs
Small neodymium cube magnet $\times 2$ BA192 \$20 Sigel
Optical adhesive NOA 61 \$30 Norland
3 mm2 CVD NV diamond DNV-B1 \$2000 Element Six
SM1 tube $\times 2$ SM1PL \$20 Thorlabs
Green LED M530L4 \$250 Thorlabs
Cornerpiece with two SM1 threads CAD file \$20 Workshop
Circuit board for diamond CAD file \$300 Eurocircuits
Hexagonal light mixing rod HMR425 \$100 Thorlabs
NBK7 right angle optical prism 45-948 \$70 Edmund Optics
638 nm longpass filter DMLP638T \$100 Thorlabs
SMA cable $\times 2$ 047-12SMPSM+ \$20 Minicircuits
Power supply for photodetector LDS12B \$70 Thorlabs
Right-angle SMA cable ACX1611-ND \$20 Digikey
SMA to BNC converter $\times 2$ 242102 \$10 Digikey
Sensitivity
@@
Example Labs
The magnetometer can be used in undergraduate labs to do quantitative analysis with ODMR spectra. For example, one can view the ODMR spectrum on the Red Pitaya user interface and alter the bias field until a [1 0 0] orientation is found [19]. This can be done by placing the two block magnets at opposite corners of the circuit board and rotating them by small amounts until two resonance peaks are seen. As a useful exercise, the strength of the bias field can be found using the equation for the frequency difference between two resonances $\Delta=2\gamma_{e}B_{z}$ after accounting for the projection factor of $0.58$ for the [1 0 0] orientation. In our case, this gives $B_{z}\approx 2.5\,\mathrm{mT}$.

### v2

@@ Abstract
preprint: APS/123-QED
I Introduction
@@
In this Note, we present a portable NV magnetometer which uses a high-power green LED instead of a laser, allowing for safe illumination of the diamond with optical powers of order 370 mW. This magnetometer prioritizes safety, portability, reliable performance, and direct visualization for outreach and undergraduate settings. The high excitation power produces fluorescence which is readily visible to the naked eye, making the system particularly effective for public demonstrations and instructional settings. Although it does not reach the affordability of the very low cost modular options which are available in the literature, the magnetometer which we present here is tailored for teaching and outreach environments where clarity of the fluorescence signal and ease and reliability of the demonstration are essential.
@@
II Concept and Demonstration
@@
The central idea of the device is to use a high-power green LED (approximately 370 mW of optical power) to excite a millimetre-scale diamond containing a high concentration of NV- centers, and to detect the resulting red photoluminescence (PL) while applying microwave radiation near the NV spin resonances. A simplified schematic of the optical and electronic layout is shown in Fig. 1 (left). The LED output is fed into a short hexagonal light-mixing rod, which homogenizes the illumination and delivers it to a right-angle prism. The prism directs the green light onto the diamond mounted on the printed circuit board, whilst simultaneously collecting the red photoluminescence and returning it through the rod to a long-pass filter and amplified photodiode. Since the green excitation and red fluorescence are both easily visible to the naked eye, students can directly observe the physical processes responsible for the ODMR signal.
@@
To clarify the optical and mechanical interface on the circuit board, Fig. 1 (upper right) provides a detailed view of the prism-rod-diamond assembly. The diamond is bonded to a reflective metal pad on the board, the prism is placed on top of the diamond, and the face of the mixing rod is aligned so that it presses against the face of the prism. This compact geometry provides efficient light collection without the need for lenses and keeps alignment requirements minimal. The board incorporates a printed single-turn loop antenna which provides the oscillating magnetic field required for ODMR. The loop is etched into the top copper layer and is driven from the SMA connector through a short copper transmission line, as shown in Fig. 1 (lower right). The diamond is positioned directly above the loop to ensure strong near-field coupling. An excerpt of the top copper layer is included in the figure for clarity, and the full Gerber files are available in the Supplementary Material.
Refer to captionRefer to caption
Figure 1: (Left) Schematic of the experiment. The gray squares denote the two block magnets attached to the circuit board and the pink square denotes the NV diamond. (Top right) Schematic showing the optical layout (not to scale). The green arrow indicates the excitation path from the LED to the diamond. The red arrow indicates NV photoluminescence emitted by the diamond and directed toward the photodetector (Bottom right) Excerpt of the top copper Gerber layer of the printed circuit board showing the integrated microwave loop used for ODMR. The circular pad on the right connects to the SMA input. A straight copper trace routes the microwave signal to the printed single-turn loop in the middle of the board, which generates the oscillating magnetic field to drive the NV spin transitions. The diamond sits directly above this loop during operation. Full PCB design files are provided in the Supplementary Material.
The Red Pitaya–based digital lock-in amplifier used in this work is implemented with open source software running on a host computer. In the present experiment, students interact with this system through a graphical user interface (GUI) that provides real-time graphical visualization of the demodulated ODMR signal, without requiring prior knowledge of digital signal processing or lock-in implementation. The interface can be operated in several ways. For ODMR spectroscopy, the microwave frequency can be swept whilst the magnitude of the lock-in response is recorded as a function of frequency, allowing ODMR resonances to be observed. Alternatively, the microwave frequency can be fixed at a selected working point and the in-phase (X) and quadrature (Y) lock-in outputs monitored as a function of time, enabling changes in the magnetic environment to be detected.
In this work, the GUI is used primarily as a qualitative visualization tool, allowing students to explore the ODMR response and its dependence on experimental parameters such as microwave frequency range, modulation depth and frequency, lock-in time constant, and LED drive current (where available). Changes to these parameters are reflected in the displayed signal, allowing students to investigate their influence on ODMR contrast, linewidth, signal-to-noise ratio, and magnetic field response. The simplified optical design eliminates the need for laser alignment, allowing students to focus on signal interpretation, data analysis, and the physics of magnetic resonance rather than optical troubleshooting.
@@
The magnetometer has been used successfully in outreach activities. The entire setup packs into two small plastic boxes and can be reassembled in a few minutes, shown in Fig. 2 (left). Fig. 2 (right) shows that both the green LED excitation and red photoluminescence can be seen with the naked eye. At a recent event in our Physics Department, students viewed the PL signal on a laptop whilst magnetic objects were brought near the sensor. For simple demonstrations, it is sufficient to observe the raw PL signal, which changes visibly in response to nearby magnetic objects. Fig. 3 shows the demodulated PL signal when a steel Allen key was moved close to the magnetometer. It can be seen that a spike forms as the object is brought near, giving a compelling demonstration that the device functions as a magnetometer.
Refer to captionRefer to caption
Figure 2: Photograph of the magnetometer during operation (left). Green LED excitation can be seen leaking from the device (top right), and red NV photoluminescence emitted by the diamond is readily observed through the long-pass filter (bottom right).
@@
III ODMR Spectroscopy
[…]
Table 1: Operating parameters used to obtain the FM ODMR spectrum shown in Fig. 2.
Parameter Value
Bias field strength $\sim 1\,\mathrm{mT}$
Modulation frequency $2.0\,\mathrm{kHz}$
Modulation amplitude $\sim 5.6\,\mathrm{MHz}$
Time constant (3 dB point) $15.29\,\mathrm{ms}$
Figure 4: Measured ODMR spectrum from the diamond magnetometer, acquired using frequency-modulation detection under the same operating conditions as the demonstration in Fig. 5. The observed resonances correspond to spin-state transitions between the Zeeman-split NV ground-state levels illustrated in Fig. S1. The plotted signal corresponds to the magnitude of the lock-in response $\sqrt{X^{2}+Y^{2}}$. The dashed line marks the working point, defined as the frequency at which the ODMR amplitude exhibits its maximum slope $dV/df$. The extrema in the FM ODMR response correspond to the steepest slopes of the underlying ODMR resonances, whose center frequencies are determined by the Zeeman splitting of the NV energy levels shown schematically in Fig. S1.
In our device, the microwave frequency is modulated and the resulting lock-in signal is therefore proportional to the derivative of the ODMR fluorescence with respect to frequency. Although this produces derivative-like lineshapes, the resonance frequencies (and hence the Zeeman splittings) are unchanged and can be directly related to the NV energy level structure. The lock-in amplification allows small changes in fluorescence to be measured with high signal-to-noise ratio. For readers unfamiliar with lock-in detection, clear pedagogical treatments can be found in [7] and in the context of NV-based ODMR measurements in Section 4D of [1].
IV Magnetometry Mode and Calibration
To verify quantitative operation of the magnetometer, the device was placed inside a Helmholtz coil and the coil current was varied whilst recording the demodulated lock-in response. For each current value, the magnitude of the lock-in signal $|R|=\sqrt{X^{2}+Y^{2}}$ was averaged over a one second interval. As shown in Fig. 5, the response varies linearly with the applied coil current over the operating range relevant for demonstrations and instructional use, confirming that the instrument provides a response which is proportional to an applied magnetic field. The lock-in output voltage is therefore used as a proxy for magnetic field strength.
Figure 5: Calibration of the magnetometer using a Helmholtz coil. The lock-in magnitude $|R|=\sqrt{X^{2}+Y^{2}}$ is plotted as a function of coil current over the linear operating range of the device, averaged over a one-second interval. The blue circles show the measured lock-in magnitude and the dashed orange line shows a linear fit, demonstrating that the response is proportional to an applied magnetic field.
@@ IV Magnetometry Mode and Calibration
V Conclusion
We have presented a portable NV center magnetometer which uses LED illumination instead of a laser, making it especially suitable for outreach and undergraduate laboratory settings. The device produces clear ODMR spectra and responds visibly to nearby magnetic objects, enabling both qualitative demonstrations and quantitative measurements. During outreach events the magnetometer proved easy to assemble and operate, with students able to observe its response in real time. In an instructional laboratory, the system can support introductory activities, as well as more advanced exercises involving magnet alignment, bias-field estimation, or parameter optimization.
We note that although most components of the magnetometer are inexpensive, the diamond represents a significant fraction of the total cost. Much cheaper alternatives based on ensembles of microdiamonds are possible and are widely used in other contexts. However, with the simple excitation and collection optics employed here, fluorescence would be collected from many randomly oriented crystallites, leading to significant broadening of the ODMR spectrum and reduced sensitivity. The present design prioritizes clarity of the spectrum, robustness, and ease of interpretation over minimal cost, which is advantageous for teaching laboratories and outreach demonstrations.
Acknowledgements.
Alex Newman’s PhD studentship was funded by an EPSRC iCASE award to UKNNL (United Kingdom National Nuclear Laboratory). This work received funding from the National Nuclear Laboratory’s Science and Technology programme (Decontamination and Decommissioning Core Science). This work is also supported by Innovate UK grant 10003146, EPSRC grant EP/V056778/1 (Prosperity Partnership with Element Six), an EPSRC Impact Acceleration Account (IAA) award, the EPSRC Q-BIOMED Hub EP/Z533191/1, the EPSRC Quantum Computation and Simulation Hub EP/T001062/1, and STFC grant number ST/W006561/1. This research is funded in part by the Gordon and Betty Moore Foundation through Grant GBMF12328, DOI 10.37807/GBMF12328. This material is based upon work supported by the Alfred P. Sloan Foundation under Grant No. G-2023-21130.
Supplementary Material: A portable LED-based diamond magnetometer for outreach and teaching labs
Hollis Williams ]holliswilliams@hotmail.co.uk Alex NewmanStuart GrahamColin StephenGavin Morley
@@ VI Background
VII LED Magnetometer
VII.1 Design
In this section, we discuss the components used to build the magnetometer (these are listed in Table I along with the model, estimated cost, and suggested supplier). The prices were checked on 29th December 2026 and there may be some variation among different vendors. It is also possible to lower the cost by looking for used parts. The overall estimated cost of the device is around \$4500, similar to the setup described in [18]. In some cases, cheaper alternatives can be used to decrease the cost (for example, filtered glass could be used instead of a dichroic shortpass filter). We also expect that some re-design of the LED part could result in a substantial price reduction.
@@ VII.1 Design
Red Pitaya STEMLab 125-14 $\mathdollar 400$ Red Pitaya
Microwave source TPI-1005-A $\mathdollar 300$ RF Consultant
Dichroic 600 nm shortpass filter 69-180 $\mathdollar 100$ Edmund Optics
Amplified photodetector PDA100A2 $\mathdollar 400$ Thorlabs
T-Cube LED driver LEDD1B $\mathdollar 300$ Thorlabs
Small neodymium cube magnet $\times 2$ BA192 $\mathdollar 20$ Sigel
Optical adhesive NOA 61 $\mathdollar 30$ Norland
33 mm2 CVD NV diamond DNV-B1 $\mathdollar 2000$ Element Six
SM1 tube $\times 2$ SM1PL $\mathdollar 20$ Thorlabs
Green LED M530L4 $\mathdollar 250$ Thorlabs
Cornerpiece with two SM1 threads CAD file $\mathdollar 20$ Workshop
Circuit board for diamond CAD file $\mathdollar 300$ Eurocircuits
Hexagonal light mixing rod HMR425 $\mathdollar 100$ Thorlabs
NBK7 right angle optical prism 45-948 $\mathdollar 70$ Edmund Optics
638 nm longpass filter DMLP638T $\mathdollar 100$ Thorlabs
SMA cable $\times 2$ 047-12SMPSM+ $\mathdollar 20$ Minicircuits
Power supply for photodetector LDS12B $\mathdollar 70$ Thorlabs
Right-angle SMA cable ACX1611-ND $\mathdollar 20$ Digikey
SMA to BNC converter $\times 2$ 242102 $\mathdollar 10$ Digikey
VII.2 Sensitivity
@@
VII.3 Sample Laboratory Exercise
This section outlines a sample laboratory exercise suitable for a 2-3 hour undergraduate teaching lab, intended to demonstrate that the magnetometer supports hands-on experimental investigation.
Learning objectives
After completing this guided laboratory exercise, students should be able to:
  • Describe the basic physical principles underlying optically detected magnetic resonance (ODMR) in NV centers.
  • Record and qualitatively interpret ODMR spectra using lock-in detection.
  • Estimate a magnetic field from measured resonance splittings.
  • Explore how experimental parameters influence signal quality and measurement uncertainty.
Experimental procedure
  1. 1.
    With no applied bias field, students record an ODMR spectrum using the Red Pitaya interface and identify the zero-field splitting.
  2. 2.
    A bias magnetic field is applied using two permanent magnets placed on the circuit board. Students adjust the magnet positions to resolve two resonance peaks corresponding to a preferred NV orientation.
  3. 3.
    Students measure the frequency separation between the resonances and estimate the magnetic field strength using the relation $\Delta=2\gamma_{e}B_{z}$, accounting for the crystallographic projection factor appropriate for the observed NV orientation.
  4. 4.
    The procedure is repeated for several magnet configurations to explore the reproducibility and uncertainty of the field estimate.
Data analysis and discussion
Students analyze how the resonance splitting depends on magnet placement and discuss the dominant sources of uncertainty, including linewidth, signal-to-noise ratio, and magnetic field alignment. The role of lock-in detection in enhancing measurement sensitivity can be explored qualitatively by varying modulation parameters within the software interface.
Extensions
This core exercise can be extended by comparing measurements with and without frequency modulation, investigating the effect of optical or microwave power on signal contrast, or introducing a calibrated external field source for quantitative comparison.
VII.4 Suggested Extensions and Project Ideas